# Identifiable Causal Inference with Noisy Treatment and No Side Information

### v2

@@ Abstract
In some causal inference scenarios, the treatment variable is measured inaccurately, for instance in epidemiology or econometrics. Failure to correct for the effect of this measurement error can lead to biased causal effect estimates. Previous research has not studied methods that address this issue from a causal viewpoint while allowing for complex nonlinear dependencies and without assuming access to side information. For such a scenario, this study proposes a model that assumes a continuous treatment variable that is inaccurately measured. Building on existing results for measurement error models, we prove that our model’s causal effect estimates are identifiable, even without knowledge of the measurement error variance or other side information. Our method relies on a deep latent variable model in which Gaussian conditionals are parameterized by neural networks, and we develop an amortized importance-weighted variational objective for training the model. Empirical results demonstrate the method’s good performance with unknown measurement error. More broadly, our work extends the range of applications in which reliable causal inference can be conducted.
@@ 1 Introduction
Causal inference deals with how a cause variable $X$, called treatment, causally affects an outcome $Y$. This is different from just estimating statistical dependencies from observational data, because even when we know that $X$ causes $Y$ and not the other way around, the statistical relationship could be affected by confounders $Z$, i.e., common causes of $X$ and $Y$. Knowing the causal relationships is crucial in fields that seek to make interventions, e.g., medicine or economics (Pearl, 2009; Peters et al., 2017; Imbens & Rubin, 2015).
@@
Causal inference may be complicated by variables being subject to noise, e.g., inaccurate measurement, clerical error, or self-reporting (often called misclassification for categorical variables). Unaccounted, it is well-known that this error can bias statistical and causal estimates in a fairly arbitrary manner, and consequently, measurement error models have been widely studied to address this bias (Carroll et al., 2006; Yi et al., 2021; Schennach, 2016; 2020). Figure 1 presents an example of this skewing effect. While it is often believed that measurement error only attenuates dependencies, this example demonstrates that on the contrary, dependencies are amplified for some ranges of the treatment variable.
Figure 1: Our method (CEME) compared against regression that assumes the data with noisy measurements of $x^{*}$ is accurate (CE-simple), and against the ground truth (GT) regression function, which is the sigmoid $S(x)=1/(1+e^{-x})$. The accurate values of $x^{*}$ are hidden from both methods, and in this example, there is no confounding. A similar example is presented by Zhu et al. (2022).
@@
To model measurement error, we use an SCM with the causal graph $G$ depicted in Figure 2. Note that the causal graph is also a Bayesian network that implies conditional independencies according to d-separation (Bishop, 2006). The variables in the graph are the confounders/covariates $Z$, the accurate treatment value $X^{*}$ which is unobserved, the noisy measurement of the treatment, denoted by $X$, and the noisy value of the outcome, denoted by $Y$. Except for $X^{*}$, the other variables are observed. The assumption of independent exogenous variables implies that there is no unobserved confounding. Therefore, when the SCM is identified, any interventional or counterfactual distributions can be computed. These include e.g. $Y_{x^{*}}$, which denotes the distribution of $Y$ after intervening with $do(X^{*}=x^{*})$ (alternatively denoted by $Y|do(X^{*}=x^{*})$), or the distribution of $Y_{x^{*}_{1}}-Y_{x^{*}_{2}}$, which compares the effects of two treatment values $x^{*}_{1}$ and $x^{*}_{2}$, or the distribution of $Y_{x^{*}}|Y=y,Z=z$, which denotes the outcome that would have been obtained from intervention $do(X^{*}=x^{*})$ when in reality no intervention was made and the values $Y=y$ and $Z=z$ were observed.
$X^{*}$ $Z$ $Y$ $X$
Figure 2: Causal graph for our model (Causal Effect estimation with Measurement Error, CEME). The observed variables $X$ (noisy treatment), $Y$ (effect/outcome), and $Z$ (confounders) are shaded to distinguish them from the hidden variable $X^{*}$ (true treatment).
@@
Because $X^{*}$ is hidden, the SCM or any causal queries that consider the effects of an intervention $do(X^{*}=x^{*})$ are not identifiable without further assumptions, as the relationship between $X^{*}$ and the other variables can not be estimated. Therefore, for identification, we rely on additional assumptions on the measurement error mechanism $X=f_{X}(X^{*},U_{X})$ such that the SCM becomes statistically identifiable. In the literature, much attention has been devoted to making the model identifiable using so called side-information, such as a known error mechanism, repeated measurements, instrumental variables, or a gold standard sample of accurate measurements (Yi et al., 2021). However, we assume no side information and instead rely on the assumptions of independence and zero-mean additive errors (Schennach & Hu, 2013): First, we assume that the measurement error is strongly classical, i.e., the observed treatment $X$ is generated by a causal mechanism
@@
where the noise $\Delta Y$ is the exogenous variable $U_{Y}$ independent of all other variables except $Y$ and has a zero mean. The conditional mean of $Y$, $\mu_{Y}(z,x^{*})={\mathbb{E}}[Y|Z=z,X^{*}=x^{*}]$, is a deterministic, possibly nonlinear function that is continuously differentiable everywhere with respect to $x^{*}$. Although not necessary for identifiability (see Section 2.3.1 for details), we further assume Gaussian distributions to facilitate practical computations, yielding
$X^{*}$ $=\mu_{X^{*}}(Z)+\sigma_{X^{*}}(Z)\cdot\Delta Z,$ (3)
$\Delta Z$ $\sim{\mathbb{N}}(0,1),$ (4)
@@
where again $\Delta Z$ is the exogenous variable $U_{Z}$. We assume that $X$, $Y$, and $X^{*}$ are real-valued scalars, although multivariate extensions are relatively straightforward. The binary (or categorical) case is briefly discussed in Appendix A. The covariate $Z$ can be a combination of categorical, discrete, or continuous variables, but the other variables are assumed to be continuous. Consequently, the learnable parameters are the variance parameters $\tau^{2}$ and $\sigma^{2}$ as well as the (deterministic) functions $\mu_{X^{*}}$, $\sigma_{X^{*}}$ and $\mu_{Y}$, which are all parameterized by neural networks to allow flexible dependencies between the variables.
We note that our model explicitly assumes measurement error only for the treatment variable $X^{*}$, and not for the covariate $Z$ or the outcome $Y$. However, measurement error in $Y$ is implicitly included in the noise variance $\sigma^{2}$ because adding Gaussian error simply increases the variance of the original Gaussian noise. To put it in another way, noise in the outcome does not bias regression estimation. On the other hand, measurement error in the covariates could be relevant, but it is restricted outside the scope of this work.
The aim of this paper is to correctly infer the SCM from observational data, given the assumptions made. To evaluate the fit, we consider the accuracy of the estimation of the following quantities: 1) the function $\mu_{Y}(z,x^{*})={\mathbb{E}}[Y|z,do(x^{*})]$, 2) the noise variances $\tau^{2}$ and $\sigma^{2}$, and 3) the function $p(y|do(x^{*}))$, which maps $x^{*}$ to the density of $Y|do(x^{*})$. The combination of 1) and 2) also evaluates the estimation accuracy of $p(y|z,do(x^{*}))$, since with $Z$ satisfying the backdoor criterion, we have $p(y|z,do(x^{*}))=p(y|z,x^{*})=\mathcal{N}(y|\mu_{Y}(z,x^{*}),\sigma^{2})$. These parameters/distributions were chosen because they are either model parameters, reflect probable use cases of the model, or both. To infer the SCM, we use deep latent variable modeling and importance weighted variational inference. Interventional distributions can then be computed on the basis of this inferred SCM. We evaluate our algorithm on a wide variety of synthetic datasets, as well as semi-synthetic data.
There is a recent, fairly rich literature on the bias introduced by measurement error in causal estimation (Yi et al., 2021). This includes qualitative analysis by encoding assumptions of the error mechanism into a causal graph (Hernán & Robins, 2021), and quantitative analysis of noise in the treatment (Zhu et al., 2022), outcome (Shu & Yi, 2019), confounders (Pearl, 2012; Miles et al., 2018) and mediators (Valeri & Vanderweele, 2014). However, most of this literature focuses on simple parametric models, such as the linear model, and the identifiability of the causal effect is achieved through side information, such as a known error mechanism, repeated measurements, instrumental variables, or a gold standard sample of accurate measurements. Exceptions include Miles et al. (2018) which studies confounder measurement error with no side information and a semiparametric model, Hu et al. (2022) which similarly to us uses deep latent variable models and variational inference, but assumes known measurement error variance, and Zhu et al. (2022) which studies treatment measurement error with a nonparametric model and notably even allowing for unobserved confounding, but utilizes both repeated measurements and an instrumental variable as side information. Schennach & Hu (2013) use non-parametric sieve estimation with no side information, without considering covariates $Z$ that could increase complexity significantly. We take them into account not only to increase the accuracy of predictions, but crucially to allow adjusting for confounders to obtain causal effects. Concurrently to us, Gao et al. (2024) study a similar model to ours and also use variational inference. However, they do not include a theoretical analysis of identifiability as we do. See Appendix A for a more broad review of identifiability results on models related to ours.
This work contributes to the literature by successfully inferring causal effects in a setting with a novel combination of challenging aspects: measurement error, no side information and complex nonlinear dependencies including dependency on confounders that must be adjusted for in order to estimate the causal effect. We also extend a previous result by Schennach & Hu (2013) to show the identifiability of our model. Finally, this work contributes by bridging different disciplines: it highlights that measurement error models constitute SCMs and uses machine learning, specifically variational inference, as a foundation for inferring causal effects.
2 Methods
2.1 Models for causal estimation
In this section, we introduce and name the models that we consider for causal estimation in our experiments:
@@ 2.1 Models for causal estimation
CEME: Causal Effect estimation with Measurement Error. This is the model defined in Section 1. The mean and standard deviation functions $\mu_{X^{*}}(Z)$, $\sigma_{X^{*}}(Z)$ and $\mu_{Y}(Z,X^{*})$ are fully connected neural networks.
@@
CE-oracle: This method assumes the model in Figure 3(a), where the true treatment $X^{*}$ is observed and is used directly for fitting the model. Hence, this model provides a loose upper bound on the performance that can be achieved by the models CEME and CEME+ which only observe the noisy treatment $X$. The model consists of only one neural network, parametrizing $\mu_{Y}(Z,X^{*})$.
CE-simple: This is the same as CE-oracle, except that it naively uses the observed noisy treatment $X$ instead of the true treatment $X^{*}$ when estimating the causal effect (model depicted in Figure 3(b)). Hence, it provides a baseline that corresponds to the usual approach of neglecting the measurement error in causal estimation. The model consists of only one neural network, parametrizing $\mu_{Y}(Z,X)$.
For inference, CEME and CEME+ use amortized variational inference to estimate the latent variable (described in the next section). In contrast, CE-oracle and CE-simple, which do not have latent variables, are trained using gradient descent with mean squared error (MSE) loss for predicting $Y$.
@@ 2.2 Inference
For the distribution $p(X^{*},X,Y|Z)$, the method requires defining a generative model (also called decoder) $p_{\theta}(x^{*},x,y|z)$, where $\theta$ denotes the parameters. In this paper, we use the model defined in Section 1 with $\theta$ consisting of the standard deviations $\tau$ and $\sigma$ as well as the functions $\mu_{X^{*}}(z)$, $\sigma_{X^{*}}(z)$ and $\mu_{Y}(z,x^{*})$, which are modeled as fully connected neural networks. (Thus in terms of the practical computations, the parameters are not the functions itself but their respective neural network weights.)
AVI also requires modeling the posterior $p(X^{*}|Z,X,Y)$ with a so-called encoder, for which we use $q_{\phi}(x^{*}|z,x,y)={\mathbb{N}}\left(x^{*}|\mu_{q}(z,x,y;\phi),\sigma_{q}(z,x,y;\phi)\right),$ where $\phi$ denotes the encoder parameters, which are the functions $\mu_{q}(z,x,y;\phi)$ and $\sigma_{q}(z,x,y;\phi)$, or alternatively, the weights of the two fully connected neural networks that used to model these functions.

### v3

@@ Abstract
In some causal inference scenarios, the treatment variable is measured inaccurately, for instance in epidemiology or econometrics. Failure to correct for the effect of this measurement error can lead to biased causal effect estimates. Previous research has not studied methods that address this issue from a causal viewpoint while allowing for complex nonlinear dependencies and without assuming access to side information. For such a scenario, this study proposes a model that assumes a continuous treatment variable that is inaccurately measured. Building on existing results for measurement error models, we prove that our model’s causal effect estimates are identifiable, even without side information and knowledge of the measurement error variance. Our method relies on a deep latent variable model in which Gaussian conditionals are parameterized by neural networks, and we develop an amortized importance-weighted variational objective for training the model. Empirical results demonstrate the method’s good performance with unknown measurement error. More broadly, our work extends the range of applications in which reliable causal inference can be conducted.
@@ 1 Introduction
Causal inference deals with how a treatment variable $X$ causally affects an outcome $Y$. This is different from just estimating statistical dependencies from observational data, because even when we know that $X$ causes $Y$ and not the other way around, the statistical relationship could be affected by confounders $Z$, i.e., common causes of $X$ and $Y$. Knowing the causal relationships is crucial in fields that seek to make interventions, e.g., medicine or economics (Pearl, 2009; Peters et al., 2017; Imbens & Rubin, 2015).
@@
Causal inference may be complicated by variables being subject to noise, e.g., inaccurate measurement, clerical error, or self-reporting (often called misclassification for categorical variables). If not accounted for, it is well-known that this error can bias statistical and causal estimates in a fairly arbitrary manner, and consequently, measurement error models have been widely studied to address this bias (Carroll et al., 2006; Buonaccorsi, 2010; Schennach, 2016; 2020).
$X^{*}$ $Z$ $Y$ $X$
Figure 1: Causal graph for our proposed model. The observed variables $X$ (noisy treatment), $Y$ (effect/outcome), and $Z$ (confounders) are shaded to distinguish them from the hidden variable $X^{*}$ (true treatment).
In this paper, we assume the treatment $X$ and outcome $Y$ to be continuous, while the confounders $Z$ can be categorical, discrete or continuous. (The case of a binary or categorical $X$ is briefly discussed in Appendix A.) While measurement error may occur in any of the variables, we explicitly model it only in $X$. We implicitly also include measurement error in $Y$, but in our model it is indistinguishable from the inherent noise in the true value of $Y$. The practical impact of this is limited, since the measurement error for $Y$ does not bias regression due to its zero-mean additive nature that we assume. On the other hand, measurement error in $Z$ could be relevant (Miles et al., 2018), but it is restricted outside the scope of this work.
The causal graph expressing our assumptions is depicted in Figure 1. An example reflecting this scenario in the context of healthcare is the effect of blood pressure on some continuous health outcome $Y$, such as arterial stiffness as measured by pulse wave velocity, or kidney function as measured by glomerular filtration rate (GFR) or by the concentration of albumin in urine. We assume $X$ is a single measurement of blood pressure, and thus it is highly subject to instantaneous variation, which plays the role of the measurement error. Some additional error probably also results from measurement apparatus inaccuracy. Here, the blood pressure that actually matters in terms of health outcomes (and whose effects we are interested in) is a longer term sliding window average, which is unobserved and which we denote by the latent variable $X^{*}$.
To enable statistical identifiability, we assume the measurement error to be additive, zero-mean and independent from the other variables. These are a standard set of assumptions in the measurement error literature, called strongly classical measurement error. Non-classical measurement error is also widely studied and well documented to occur in practice (see Schennach (2016) for a review), but it is restricted outside the scope of this work. However, we believe our modeling and inference methodology is general enough that extensions to these cases are fairly straightforward. Figure 2 presents a synthetic example of the skewing effect of measurement error in the treatment. While it is often believed that this type of measurement error only introduces attenuating bias to a naive estimate not accounting for the error (Yi et al., 2021), this example demonstrates that on the contrary, amplification is also possible. Here attenuation happens only around $X^{*}=0$, but for small and large $X^{*}$ the naive estimate amplifies the strength of the dependency.
Figure 2: Comparison of our method (CEME) against a naive method which does not account for measurement error in the treatment $X^{*}$. The accurate values of $X^{*}$ are hidden from both methods. Ground truth is the true mean function of the data generating process. The same data are displayed both with and without measurement error in $X^{*}$. It can be seen that our method “CEME” fits the error-free data (even if they are not seen by any method) whereas the “naive” method fits the data with error and cannot estimate the true regression function accurately. A similar example is presented by Zhu et al. (2022).
The measurement error problem is often addressed by making additional measurements to enable model identification. The information obtained in this way could be a known measurement error variance, or so-called side information, such as repeated measurements, instrumental variables, or a gold-standard sample of accurate measurements (Yi et al., 2021). However, we study the scenario where none of this is available, and we must rely on assumptions that could reasonably be made a priori, such as the error being additive and having a zero mean. Further, we study the scenario where the dependencies between the variables can be complex and nonlinear, and an observed confounder is present. There is a gap in the literature for causal inference when all of these challenging aspects are present.
To address this gap, our paper provides a solution by inferring a structural causal model (SCM) from observational data using deep latent variable modeling and importance weighted variational inference (Zhang et al., 2019). The inferred SCM enables the computation of any interventional and counterfactual distributions on the variables in the model, but to evaluate the fit, we consider the accuracy of the estimation of the following quantities: 1) the function $\mu_{Y}(z,x^{*})={\mathbb{E}}[Y|z,do(x^{*})]$, 2) the noise variances $\tau^{2}$ and $\sigma^{2}$, and 3) the function $p(y|do(x^{*}))$, which maps $x^{*}$ to the density of $Y|do(x^{*})$. The combination of 1) and 2) also evaluates the estimation accuracy of $p(y|z,do(x^{*}))$, since with $Z$ satisfying the backdoor criterion and by assuming a conditionally Gaussian $Y$, we have $p(y|z,do(x^{*}))=p(y|z,x^{*})=\mathcal{N}(y|\mu_{Y}(z,x^{*}),\sigma^{2})$. These parameters/distributions were chosen because they are either model parameters, reflect probable use cases of the model, or both. To facilitate practical computation, we assume conditionally Gaussian variables. However, we emphasize that the purpose of this is to simplify computations, and the identification of the model does not rely on this assumption, see Section 2.3.1 for the details. We leave up to future work to relax it by using flexible probability density estimators such as normalizing flows. We evaluate our algorithm on a wide variety of synthetic datasets, as well as semi-synthetic data.
There is previous literature on all of the challenges tackled in this work, although not previously studied all together. For an overview of the literature on the bias introduced by measurement error in causal estimation, see Yi et al. (2021). This literature includes qualitative analysis by encoding assumptions of the error mechanism into a causal graph (Hernán & Robins, 2021), and quantitative analysis of noise in the treatment (Zhu et al., 2022), outcome (Shu & Yi, 2019), confounders (Pearl, 2012; Miles et al., 2018) and mediators (Valeri & Vanderweele, 2014). However, most of this literature focuses on simple parametric models, such as the linear model, and the identifiability of the causal effect is achieved through a fully known error mechanism or side information, such as repeated measurements, instrumental variables, or a gold standard sample of accurate measurements. Exceptions include Miles et al. (2018) which studies confounder measurement error with no side information and a semiparametric model, Hu et al. (2022) which similarly to us uses deep latent variable models and variational inference, but assumes known measurement error variance, and Zhu et al. (2022) which studies treatment measurement error with a nonparametric model and notably even allowing for unobserved confounding, but utilizes both repeated measurements and an instrumental variable as side information. Schennach & Hu (2013) use non-parametric sieve estimation with no side information, without considering covariates $Z$ that could increase complexity significantly. We take them into account not only to increase the accuracy of predictions, but crucially to allow adjusting for confounders to obtain causal effects. Concurrently to us, Gao et al. (2024) study a similar model to ours and also use variational inference. However, they do not include a theoretical analysis of identifiability as we do. See Appendix A for additional references to identifiability results on models related to ours.
To summarize, this work contributes to the literature by successfully inferring causal effects in a setting with a novel combination of challenging aspects: measurement error, no side information, no knowledge of the measurement error variance, and complex nonlinear dependencies including dependency on confounders that must be adjusted for in order to estimate the causal effect. To show the identifiability of our model, we extend a previous result by Schennach & Hu (2013) to allow for confounders/covariates, and adjust it to the specific assumptions of our model. Finally, this work contributes by bridging different disciplines: it highlights that measurement error models constitute SCMs and uses machine learning, specifically variational inference, as a foundation for inferring causal effects.
2 Methods
2.1 Models for causal estimation
@@
To model measurement error, we use an SCM with the causal graph $G$ depicted in Figure 1. Note that the causal graph is also a Bayesian network that implies conditional independencies according to d-separation (Bishop, 2006). The variables in the graph are the confounders/covariates $Z$, the accurate treatment value $X^{*}$ which is unobserved, the noisy measurement of the treatment, denoted by $X$, and the noisy value of the outcome, denoted by $Y$. Except for $X^{*}$, the other variables are observed. The SCM with independent noise terms and no hidden confounders implies the common assumption of causal sufficiency. Therefore, when the SCM is identified, any interventional or counterfactual distributions can be computed. These include e.g. $Y_{x^{*}}$, which denotes the distribution of $Y$ after intervening with $do(X^{*}=x^{*})$ (alternatively denoted by $Y|do(X^{*}=x^{*})$), or the distribution of $Y_{x^{*}_{1}}-Y_{x^{*}_{2}}$, which compares the effects of two treatment values $x^{*}_{1}$ and $x^{*}_{2}$, or the distribution of $Y_{x^{*}}|Y=y,Z=z$, which denotes the outcome that would have been obtained from intervention $do(X^{*}=x^{*})$ when in reality no intervention was made and the values $Y=y$ and $Z=z$ were observed.
@@
Because $X^{*}$ is hidden, the SCM or any causal queries that consider the effects of an intervention $do(X^{*}=x^{*})$ are not identifiable without further assumptions, as the relationship between $X^{*}$ and the other variables can not be estimated. Therefore, for identification, we rely on additional assumptions on the measurement error mechanism $X=f_{X}(X^{*},U_{X})$ such that the SCM becomes statistically identifiable. In the literature, much attention has been devoted to making the model identifiable using a known error variance or side information. However, we instead rely on the assumptions of independence and zero-mean additive errors (Schennach & Hu, 2013): First, we assume that the measurement error is strongly classical, i.e., the observed treatment $X$ is generated by a causal mechanism
@@
where the noise $\Delta Y$ is the exogenous variable $U_{Y}$ independent of all other variables except $Y$ and has a zero mean. The conditional mean of $Y$, $\mu_{Y}(z,x^{*})={\mathbb{E}}[Y|Z=z,X^{*}=x^{*}]$, is a deterministic, possibly nonlinear function that is continuously differentiable everywhere with respect to $x^{*}$. (Note that by uppercase letters we denote random variables and by lowercase letters their specific realized values. Thus $\mu_{Y}(z,x^{*})$ refers to the value of $\mu_{Y}$ with argument values $z$ and $x^{*}$, while $\mu_{Y}(Z,X^{*})$ denotes the random variable that is obtained by applying $\mu_{Y}$ to the random variables $Z$ and $X^{*}$.) Although not necessary for identifiability (see Section 2.3.1 for details), we further assume Gaussian distributions to facilitate practical computations, yielding
$X^{*}|Z$ $\sim{\mathbb{N}}(\mu_{X^{*}}(Z),\sigma^{2}_{X^{*}}(Z))$ (3)
@@
We assume that $X$, $Y$, and $X^{*}$ are real-valued scalars, although multivariate extensions are relatively straightforward. The covariate $Z$ can be a combination of categorical, discrete, or continuous variables, but the other variables are assumed to be continuous. Consequently, the learnable parameters are the variance parameters $\tau^{2}$ and $\sigma^{2}$ as well as the (deterministic) functions $\mu_{X^{*}}$, $\sigma_{X^{*}}$ and $\mu_{Y}$, which are all parameterized by neural networks to allow flexible dependencies between the variables. The zero-mean assumptions in Equations (4) and (5) are made to achieve statistical identifiability. Otherwise, for any non-zero mean $\mu_{\Delta Y}$, we could obtain a new observationally equivalent model by using $\mu^{\prime}_{Y}(z,x^{*})=\mu_{Y}(z,x^{*})-\mu_{\Delta Y}$. Similarly, for any non-zero mean $\mu_{\Delta X}$, we could obtain a new observationally equivalent model by using $\mu^{\prime}_{X^{*}}(z)=\mu_{X^{*}}(z)-\mu_{\Delta X}$ and $\mu^{\prime}_{Y}(z,x^{*})=\mu_{Y}(z,x^{*}+\mu_{\Delta X})$.
Finally, we define the following model variants that we evaluate against each other:
@@
CEME: Causal Effect estimation with Measurement Error. This is the model defined in Equations (1)–(5). The mean and standard deviation functions $\mu_{X^{*}}(Z)$, $\sigma_{X^{*}}(Z)$ and $\mu_{Y}(Z,X^{*})$ are fully connected neural networks.
@@
Oracle: This method assumes the model in Figure 3(a), where the true treatment $X^{*}$ is observed and is used directly for fitting the model. Hence, this model provides a loose upper bound on the performance that can be achieved by the models CEME and CEME+ which only observe the noisy treatment $X$. The model consists of only one neural network, parametrizing $\mu_{Y}(Z,X^{*})$.
Naive: This is the same as Oracle, except that it naively uses the observed noisy treatment $X$ instead of the true treatment $X^{*}$ when estimating the causal effect (model depicted in Figure 3(b)). Hence, it provides a baseline that corresponds to the usual approach of neglecting the measurement error in causal estimation. The model consists of only one neural network, parametrizing $\mu_{Y}(Z,X)$.
To summarize the roles of the methods, CEME is our proposed method, CEME+ enables comparison with the case of a known measurement error variance $\tau^{2}$, Oracle is a loose upper bound for performance, and Naive is a baseline. For inference, CEME and CEME+ use amortized variational inference to estimate the latent variable (described in the next section). In contrast, Oracle and Naive, which do not have latent variables, are trained using gradient descent with mean squared error (MSE) loss for predicting $Y$.
@@ 2.2 Inference
For the distribution $p(X^{*},X,Y|Z)$, the method requires defining a generative model (also called decoder) $p_{\theta}(x^{*},x,y|z)$, where $\theta$ denotes the parameters. In this paper, we use the CEME model defined in Equations (1)–(5) with $\theta$ consisting of the standard deviations $\tau$ and $\sigma$ as well as the functions $\mu_{X^{*}}(z)$, $\sigma_{X^{*}}(z)$ and $\mu_{Y}(z,x^{*})$, which are modeled as fully connected neural networks. (Thus in terms of the practical computations, the parameters are not the functions itself but their respective neural network weights.)
AVI also requires modeling the posterior $p(X^{*}|Z,X,Y)$ with a so-called encoder, for which we use $q_{\phi}(x^{*}|z,x,y)={\mathbb{N}}\left(x^{*}|\mu_{q}(z,x,y;\phi),\sigma_{q}(z,x,y;\phi)\right),$ where $\phi$ denotes the encoder parameters, which are the functions $\mu_{q}(z,x,y;\phi)$ and $\sigma_{q}(z,x,y;\phi)$, or alternatively, the weights of the two fully connected neural networks that model these functions.